Check list to go from Controlled to Restricted Access

October 25, 2008

           ~cdaq/gen/Documents/check_to_restricted.tex

Date: Time: Shift Leader: Run Coordinator: Reason to open hall:

To change the hall from Controlled Access to Restricted Access, the following sequence of events must happen (please initial item as it is completed):

Initials

• Shift leader (or Run Coordinator if no shifts are underway) determines that the hall must change to Restricted Access.

• Shift leader has verbal concurrence from the Run Coordinator to change to Restricted Access.

• Shift leader has verbal concurrence from either the Hall C Work Coordinator or the Physics Liaison to change to Restricted Access.

• Run Coordinator or the Shift leader requests the Target Coordinator (or designee) to ramp down the target field.

• Field is down.

• Hall C Work Coordinator (or designee) and an ARM visually inspect the target chamber window coverings.

• Complete radiological survey.

• Target Operator disconnects the magnet leads.

• Hall C Work Coordinator locks leads in box.

• Shift Leader requests MCC to change the state of the Hall to Restricted Access.

• Hall C Work Coordinator (or designee) must supervise activities that necessitated the change to Restricted Access.

• Revert to Controlled Access as soon as possible.